# IndoNLP 2025: Shared Task on Real-Time Reverse Transliteration for Romanized Indo-Aryan languages

### v2

@@ 2 Related Works
Recent advancements in deep learning have revolutionized transliteration tasks. Neural network-based sequence-to-sequence models have gained traction, particularly those utilizing recurrent architectures. Kunchukuttan et al. (2015) introduced Brahmi-Net, a transliteration system leveraging statistical approaches for 18 Indo-Aryan languages. Nanayakkara et al. (2022) employed a Bidirectional LSTM and LSTM for Sinhala transliteration, incorporating character-level context. Similarly, Baiju et al. (2024) utilized an attention-based Bi-LSTM for Malayalam transliteration. Transformer architectures have emerged as state-of-the-art for transliteration due to their ability to handle long-range dependencies with self-attention mechanisms. Zohrabi et al. (2023) employed Transformer-based models for Azerbaijani transliteration while Madhani et al. (2023) introduced IndicXlit, a multilingual system leveraging Transformer architecture. Pretrained multilingual models like mT5 and M2M100 have also been fine-tuned for transliteration tasks, treating transliteration as a translation problem. A deep learning approach proposed by Deselaers et al. (2009) used a deep belief network for Arabic-English transliteration, and a few other Arabizi to Arabic transliteration are explored Masmoudi et al. (2019); Al-Badrashiny et al. (2014).
@@ 3.5 Participant’s Systems
The IndiDataMiner team at the Indian Institute of Technology Guwahati proposed a sentence-level back-transliteration approach using the LLaMa 3.1 model for Hindi. Their approach addresses the challenges posed by the increasing use of Romanized typing for Indo-Aryan languages on social media, which often lacks standardization and results in loss of linguistic richness. To resolve the ambiguities in Romanized Hindi text, they leveraged fine-tuning with the Dakshina dataset. The team’s approach includes both zero-shot learning and fine-tuning techniques to enhance the model’s transliteration capabilities. The model was trained using the Dakshina dataset, which provides a parallel corpus of 12 Indian languages, including Hindi. The dataset was formatted to fit the Alpaca prompt structure, with the instruction to transliterate Romanized Hindi back to Devanagari script, ensuring consistency across training and testing. They used the LLaMA 3.1 8B model, a large-scale transformer-based architecture, which was optimized for causal language modeling and enhanced with Low-Rank Adaptation (LoRA) and 4-bit quantization. The training process utilized the SFTTrainer class from the trl library and the Unsloth framework for 4-bit quantization, which improved memory efficiency. The fine-tuned model demonstrated significant improvements in transliteration accuracy, achieving high BLEU scores on the Hindi test dataset.
@@
Team Vectora’s approach to Sinhala back-transliteration is a context-aware system that combines multiple techniques to handle the complexities of "Singlish". Their method employs several sophisticated components that work together to achieve accurate transliteration. At the core of their system is a dictionary-based mapping approach that uses an ad-hoc transliteration dictionary to map common Singlish words to their Sinhala equivalents. This dictionary is particularly effective at handling frequent words and their different possible Sinhala representations. To accommodate the complexity of the language, the dictionary incorporates multiple mappings for ambiguous words. For handling words not found in the dictionary (out-of-vocabulary words), the system implements rule-based transliteration based on Sinhala phonetic patterns. This component proves essential for processing words that may not be common or are absent from their dictionary. To resolve lexical ambiguities where one Romanized word can map to multiple Sinhala words, the system employs a BERT-based language model to analyze the sentence-level context and select the most appropriate Sinhala word. The contextual disambiguation process works through a sophisticated sequence of operations. The system first generates all possible sentences by filling the masked ambiguous words with candidate Sinhala words. These generated sentences are then evaluated using a BERT model configured for Masked Language Modeling (MLM). The final transliterated output is determined by selecting the sentence that achieves the highest probability score. To address processing time challenges, the system incorporates several optimization strategies. These include reducing candidate words for ambiguous words through a filtering mechanism and implementing sentence chunking based on the number of BERT calls. The effectiveness of the entire system is rigorously evaluated using multiple metrics.
@@
The Team MoreCSE explored two distinct approaches for Sinhala transliteration: a rule-based system and a deep learning-based system. The rule-based approach implements a systematic method using predefined linguistic rules to map Latin script (Singlish) to Sinhala script. This system operates through a character-by-character matching strategy, processing each input word by matching the longest possible substring (up to three characters) with rules in a transliteration table. The system follows a straightforward logic: when a match is found, the corresponding Sinhala character is appended to the result; if no match is found, the character is added unchanged. This method builds upon and enhances a previous rule-based system by incorporating additional rules for two and three-character mappings. Their deep learning approach takes a fundamentally different path by modelling transliteration as a translation task, leveraging a Transformer-based encoder-decoder architecture. The team fine-tuned the M2M100 model, a pre-trained multilingual sequence-to-sequence model, by treating Romanized Sinhala as the English input and the Sinhala script as the target language. This implementation utilizes the M2M100’s tokenizer and learns the intricate relationships between the two language pairs. The deep learning method offers several key advantages: it enables context-based generation, eliminates the need for manual rule definition, and can effectively handle code-mixed and code-switched cases through expanded training data. The model’s training was conducted on a substantial dataset comprising 10k parallel data points. Their analysis revealed that while the deep learning approach demonstrated superior robustness in handling language variability, it proved less computationally efficient than its rule-based counterpart.
@@
The approach of Team NexText for Malayalam transliteration utilizes Bi-LSTM layers. They trained their model on a substantial combined dataset of 4.3 million transliteration pairs from the Dakshina and Aksharantar datasets. Their data preprocessing strategy operates at the word level, with the model trained on transliteration pairs. During testing, sentences undergo preprocessing, where individual words are extracted, transliterated independently, and then reconstructed into complete sentences. The system carefully preserves non-alphabetic characters, reinserting them after the complete transliteration process. The model architecture is comprehensive and multi-layered. It begins with an encoder input layer capable of processing up to 57 characters in length character sequences. These characters are transformed into 64-dimensional vectors through an embedding layer. A bidirectional LSTM layer then processes this information, capturing sequence patterns from both directions and creating a 256-dimensional representation. This representation is refined through a dense layer, which reduces the dimensionality to a 128-dimensional vector, creating a context vector that feeds into the decoder. The decoder architecture employs a repeat vector layer to duplicate the context vector for each timestep, followed by an LSTM layer for generating hidden states. An attention mechanism enhances the model’s ability to focus on relevant parts of the input sequence during the decoding process. The final stage combines the LSTM decoder output with the attention layer through concatenation, feeding into a time-distributed dense layer that produces a probability distribution over possible output characters. The training was conducted on a high-performance computing setup, specifically a single Nvidia DGX A100 GPU with 80 GB RAM. The evaluation was conducted on two distinct test sets: one focusing on general transliteration patterns and another featuring ad-hoc patterns with frequent vowel omissions. While the model demonstrated strong performance on standard typing patterns, it showed some performance degradation when handling ad-hoc typing patterns.

### v3

@@ 2 Related Works
Recent advancements in deep learning have revolutionized transliteration tasks. Neural network-based sequence-to-sequence models have gained traction, particularly those utilizing recurrent architectures. Kunchukuttan et al. (2015) introduced Brahmi-Net, a transliteration system leveraging statistical approaches for 18 Indo-Aryan languages. Nanayakkara et al. (2022) employed a Bidirectional LSTM and LSTM for Sinhala transliteration, incorporating character-level context. Transformer architectures have emerged as state-of-the-art for transliteration due to their ability to handle long-range dependencies with self-attention mechanisms. Zohrabi et al. (2023) employed Transformer-based models for Azerbaijani transliteration while Madhani et al. (2023) introduced IndicXlit, a multilingual system leveraging Transformer architecture. Pretrained multilingual models like mT5 and M2M100 have also been fine-tuned for transliteration tasks, treating transliteration as a translation problem. A deep learning approach proposed by Deselaers et al. (2009) used a deep belief network for Arabic-English transliteration, and a few other Arabizi to Arabic transliteration are explored Masmoudi et al. (2019); Al-Badrashiny et al. (2014).
@@ 3.5 Participant’s Systems
The IndiDataMiner team Kumar et al. (2025) at the Indian Institute of Technology Guwahati proposed a sentence-level back-transliteration approach using the LLaMa 3.1 model for Hindi. Their approach addresses the challenges posed by the increasing use of Romanized typing for Indo-Aryan languages on social media, which often lacks standardization and results in loss of linguistic richness. To resolve the ambiguities in Romanized Hindi text, they leveraged fine-tuning with the Dakshina dataset. The team’s approach includes both zero-shot learning and fine-tuning techniques to enhance the model’s transliteration capabilities. The model was trained using the Dakshina dataset, which provides a parallel corpus of 12 Indian languages, including Hindi. The dataset was formatted to fit the Alpaca prompt structure, with the instruction to transliterate Romanized Hindi back to Devanagari script, ensuring consistency across training and testing. They used the LLaMA 3.1 8B model, a large-scale transformer-based architecture, which was optimized for causal language modeling and enhanced with Low-Rank Adaptation (LoRA) and 4-bit quantization. The training process utilized the SFTTrainer class from the trl library and the Unsloth framework for 4-bit quantization, which improved memory efficiency. The fine-tuned model demonstrated significant improvements in transliteration accuracy, achieving high BLEU scores on the Hindi test dataset.
@@
Team Vectora’s Perera et al. (2025) approach to Sinhala back-transliteration is a context-aware system that combines multiple techniques to handle the complexities of "Singlish". Their method employs several sophisticated components that work together to achieve accurate transliteration. At the core of their system is a dictionary-based mapping approach that uses an ad-hoc transliteration dictionary to map common Singlish words to their Sinhala equivalents. This dictionary is particularly effective at handling frequent words and their different possible Sinhala representations. To accommodate the complexity of the language, the dictionary incorporates multiple mappings for ambiguous words. For handling words not found in the dictionary (out-of-vocabulary words), the system implements rule-based transliteration based on Sinhala phonetic patterns. This component proves essential for processing words that may not be common or are absent from their dictionary. To resolve lexical ambiguities where one Romanized word can map to multiple Sinhala words, the system employs a BERT-based language model to analyze the sentence-level context and select the most appropriate Sinhala word. The contextual disambiguation process works through a sophisticated sequence of operations. The system first generates all possible sentences by filling the masked ambiguous words with candidate Sinhala words. These generated sentences are then evaluated using a BERT model configured for Masked Language Modeling (MLM). The final transliterated output is determined by selecting the sentence that achieves the highest probability score. To address processing time challenges, the system incorporates several optimization strategies. These include reducing candidate words for ambiguous words through a filtering mechanism and implementing sentence chunking based on the number of BERT calls. The effectiveness of the entire system is rigorously evaluated using multiple metrics.
@@
The Team MoreCSE De Mel et al. (2024) explored two distinct approaches for Sinhala transliteration: a rule-based system and a deep learning-based system. The rule-based approach implements a systematic method using predefined linguistic rules to map Latin script (Singlish) to Sinhala script. This system operates through a character-by-character matching strategy, processing each input word by matching the longest possible substring (up to three characters) with rules in a transliteration table. The system follows a straightforward logic: when a match is found, the corresponding Sinhala character is appended to the result; if no match is found, the character is added unchanged. This method builds upon and enhances a previous rule-based system by incorporating additional rules for two and three-character mappings. Their deep learning approach takes a fundamentally different path by modelling transliteration as a translation task, leveraging a Transformer-based encoder-decoder architecture. The team fine-tuned the M2M100 model, a pre-trained multilingual sequence-to-sequence model, by treating Romanized Sinhala as the English input and the Sinhala script as the target language. This implementation utilizes the M2M100’s tokenizer and learns the intricate relationships between the two language pairs. The deep learning method offers several key advantages: it enables context-based generation, eliminates the need for manual rule definition, and can effectively handle code-mixed and code-switched cases through expanded training data. The model’s training was conducted on a substantial dataset comprising 10k parallel data points. Their analysis revealed that while the deep learning approach demonstrated superior robustness in handling language variability, it proved less computationally efficient than its rule-based counterpart.
@@
The approach of Team NexText Baiju et al. (2024) for Malayalam transliteration utilizes Bi-LSTM layers. They trained their model on a substantial combined dataset of 4.3 million transliteration pairs from the Dakshina and Aksharantar datasets. Their data preprocessing strategy operates at the word level, with the model trained on transliteration pairs. During testing, sentences undergo preprocessing, where individual words are extracted, transliterated independently, and then reconstructed into complete sentences. The system carefully preserves non-alphabetic characters, reinserting them after the complete transliteration process. The model architecture is comprehensive and multi-layered. It begins with an encoder input layer capable of processing up to 57 characters in length character sequences. These characters are transformed into 64-dimensional vectors through an embedding layer. A bidirectional LSTM layer then processes this information, capturing sequence patterns from both directions and creating a 256-dimensional representation. This representation is refined through a dense layer, which reduces the dimensionality to a 128-dimensional vector, creating a context vector that feeds into the decoder. The decoder architecture employs a repeat vector layer to duplicate the context vector for each timestep, followed by an LSTM layer for generating hidden states. An attention mechanism enhances the model’s ability to focus on relevant parts of the input sequence during the decoding process. The final stage combines the LSTM decoder output with the attention layer through concatenation, feeding into a time-distributed dense layer that produces a probability distribution over possible output characters. The training was conducted on a high-performance computing setup, specifically a single Nvidia DGX A100 GPU with 80 GB RAM. The evaluation was conducted on two distinct test sets: one focusing on general transliteration patterns and another featuring ad-hoc patterns with frequent vowel omissions. While the model demonstrated strong performance on standard typing patterns, it showed some performance degradation when handling ad-hoc typing patterns.